OpenStudy (anonymous):

Please Help!! Find the rectangular coordinates of the point with the polar coordinates (1, 1/2 pi).

OpenStudy (anonymous):

@phi

OpenStudy (phi):

use x= r cos(A) and y = r sin(A) (which if you know trig, you could have puzzled out by noticing you have a triangle and you need to find its base (x value) and height (y value)

OpenStudy (anonymous):

I'm in pre-cal but I don't understand any of this

OpenStudy (anonymous):

1=r cos(A) 1 1/2 pi=r sin(A)?

OpenStudy (phi):

Let's start by getting the ideas straight. when they write (1, 1/2 pi) they mean that pair to stand for (r, A) where r is the distance from the origin, and A is the direction (the angle) Unfortunately, the way they write it looks just like (x,y) In fact, if you were not told the polar coordinates (1, 1/2 pi). there would be no way to know you did NOT have (x,y) coordinates.

OpenStudy (anonymous):

ok, so x=1 cos(1 1/2pi) y=1 sin(1 1/2pi)

OpenStudy (phi):

yes, but what is 1 ½ pi ? the angle is the second number in the pair (1, ½ pi) A is pi/2

OpenStudy (anonymous):

oh gosh, i kept reading that wrong sorry

OpenStudy (phi):

When you have time, watch this (it will probably help a lot) http://www.khanacademy.org/math/precalculus/parametric_equations/polar_coor/v/polar-coordinates-1

OpenStudy (anonymous):

ok

OpenStudy (phi):

so you should get x=0 y=1

OpenStudy (anonymous):

i kept getting x=1 y=0

OpenStudy (phi):

x= cos(pi/2) y = sin(pi/2) pi/2 is the same as 90º

OpenStudy (phi):

Did you use a calculator? Are you in radian mode? btw, people memorize the sin, cos and tan of pi/2

OpenStudy (anonymous):

ohh ok i'm not in radian mode

OpenStudy (anonymous):

Thank you :)

OpenStudy (phi):

yw